OpenStudy (precal):

question about calculus notation slope at a point is f'(a) what is d/dx (f(lnx))?

OpenStudy (precal):

Isn't the second one to take the derivative of the composition function f(lnx)?

OpenStudy (anonymous):

Chain rule applies \[\frac{ d(f(\ln x))}{ dx } = f \prime(\ln x) \times \frac{f(\ln x) }{ x }\]

OpenStudy (precal):

ok I will take another look at it but I was working from inside out ie I sub lnx then took the derivative of the function

OpenStudy (anonymous):

Maybe the reason the problem mentions \[f \prime (a)\] is to let you know that f(x) is a differentiable function. Therefore \[f \prime (\ln (x))\] exists.

OpenStudy (precal):

no actually f'(a) is not mentioned at all. Someone I was tutoring did not understand the difference in notation. I just wanted another point of view.